The mean of 10 scores is 14. Another 2 scores are included and the new mean is 16. What is the mean of the two new scores?
Answer by stanbon(57984)

The mean of 10 scores is 14.
---
Sum of the 10 scores is 10*14 = 140
----------------------
Another 2 scores are included and the new mean is 16.
Sum of the 12 scores is 12*16 = 192
-----------------------
What is the mean of the two new scores?
Sum of the scores of the 2 added scores = 192-140 = 52
-----------
Average of the 2 added scores = 52/2 = 26
===============================================
Cheers,
Stan H.
============